# Rethinking Multi-modal Object Detection from the Perspective of Mono-Modality Feature Learning

### v1

@@ 4.1 Datasets and Evaluation Metrics
1) DroneVehicle: This dataset [27] includes images captured by drones in urban areas under different lighting conditions. It has annotations with oriented bounding boxes for five categories: ‘car‘, ‘truck’, ‘bus’, ‘van’, and ‘freight car’. There are 28,439 pairs of RGB and IR images, with 17,990 for training, 1,469 for validation, and 8,980 for test.
@@
4) Mean Average Precision (mAP): mAP is a standard metric for object detection. It evaluates detection performance based on classification accuracy and overlap (IoU). mAP50 is the average precision at an IoU of 0.50 and mAP averages mAP across IoU thresholds from 0.50 to 0.95.
@@ 4.2 Implementation Details
All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs. We employs CSPDarknet53 as the backbone network, and modified it to process dual-modal inputs. The network is optimized using SGD with a momentum of $0.937$ and a weight decay of $5\times 10^{-4}$. We initialize the learning rate at $1\times 10^{-2}$ and gradually decrease it to $1\times 10^{-4}$ throughout the training process. In the testing phase, we set the confidence threshold to $0.25$ to filter detection results. If there is no special state, the ablation study results on the DroneVehicle dataset are conducted on the validation set.
@@
Figure 5: Visualization of the Fusion Degradation phenomenon across the DroneVehicle (first and second rows) and FLIR (third and fourth rows) datasets. Columns from left to right represent infrared modality, RGB modality, naive add method, weighted-based method, attention-based method, our M2D-LIF and the ground truth. The Green boxes indicate correctly detected objects, while the red dashed boxes represent the Fusion Degradation phenomenon, where objects detected by mono-modal models but missed by the multi-modal object detection model. The Orange boxes emphasize how our proposed method effectively resolves this issue. Zoom in for details.
Figure 6: Visualization of the Weight Map $W_{V}$ in Equation 8. From left to right are the infrared image, RGB image, the weight map and the grayscale color corresponding to the value.
4.3 Ablation Study
1) Study on Each Component: To evaluate the effectiveness of the M2D-LIF framework, we conducted ablation studies on the M2D and LIF modules. We incrementally applied these components to the baseline model to evaluate their individual contributions, as shown in Table 2. Without any of these two modules, the model achieves mAP scores of 44.4%, 66.9%, and 68.2% on the FLIR, DroneVehicle, and LLVIP datasets, respectively. The integration of the M2D training improves model performance by 1.6% in FLIR, 2.3% in DroneVehicle, and 2.5% in LLVIP. The best results are obtained by combining both the M2D and LIF modules, resulting in increases of 2.6%, 3.7%, and 2.6% over the baseline model. These results highlight the impact the significant impact of the complete M2D-LIF framework.
2) Study on the LIF: To demonstrate that our proposed LIF module surpasses simply sensing illumination, we conducted a comparative experiment against the method directly using the normalized L-Channel from LAB color space as fusion weights. As shown in Table 3, our LIF module outperforms this approach by 1.9%, 3.2%, and 0.6% on the FLIR, DroneVehicle, and LLVIP datasets, respectively. Simply using the normalized L-Channel as the fusion weight cannot enable dynamic learning of the fusion weight, which is not conducive to the complementary fusion of multiple modalities. In the FLIR dataset, it even underperforms compared to the baseline method. These results further confirm that our LIF module can not only utilize the illumination context but also effectively leverage the complementary relationship between different modalities.
3) Study on the Hyper-parameter $\beta$: To investigate the impact of the $\beta$ value in the Equation 8 on M2D-LIF framework performance, we conducted ablation studies across a pre-defined set of $\beta$ values. As shown in Figure 4, The bar chart represents the AP of different classes and the line chart represents mAP. When beta takes the value of 0.4, our method achieves the best result. If the value of beta exceeds 0.8, it may lead to the modality imbalance, thus causing a serious decline in performance. More ablation studies are detailed in the supplementary material.
@@ 4.4 Visualization
Table 4: Camparison of the performance measured by mAP50, mAP on the DroneVehicle dataset. We replace the ’HBB’ detectors to the ’OBB’ detectors. The best results are highlighted in red and the second-place are highlighted in blue.
@@ 4.5 Comparison With State-of-the-Art Methods
1) On DroneVehicle: We evaluated the detection performance and parameter count of our method against four mono-modal oriented bounding box (OBB) detection methods, as well as eight SOTA multi-modal OBB detection methods on the DroneVehicle test and validation set. Table 4 presents the comparative results. Notably, our M2D-LIF framework achieves the highest mAP50 and mAP of 81.4% and 68.1% on the test set. Additionally, it recorded the highest mAP50 of 85.2% on the validation set. Furthermore, our method attained the highest mAP50 across all five categories on the validation set and three first-place and two second-place rankings on the test set. Our method has not only achieved state-of-the-art (SOTA) performance, but also the parameter count of our M2D-LIF framework is the optimal with 37.1M. It has only increased by 0.1M compared to the baseline model in the test phase, which demonstrates its superiority over other multi-modal methods.
2) On FLIR and LLVIP: Furthermore, we conducted comprehensive comparisons of our method against state-of-the-art object detection methods on the FLIR and LLVIP datasets, involving four mono-modal and seven multi-modal methods, as presented in Table 5. Our M2D-LIF framework achieves 46.1% and 70.8% of mAP on the FLIR and LLVIP datasets respectively, while only employing 36.6M parameters, the second lowest among multi-modal methods. The above results underscore the effectiveness of our M2D-LIF framework in handling diverse datasets. The superior performance on all the three datasets highlights the adaptability of the proposed framework. Our method not only competes with but often surpasses existing state-of-the-art multi-modal methods while achieving higher computational efficiency with a lower parameter.
@@ 5 Conclusions
In this paper, we first observed the fusion degradation phenomenon. Through linear probing evaluation, we further identified mono-modality insufficient learning as its underlying cause of this phenomenon. To address this issue, we rethought the multi-modal object detection task from the mono-modality learning perspective and constructed an end-to-end M2D-LIF framework. The Mono-Modality Distillation (M2D) method was designed to enhance enhancing the feature extraction capability of the multi-modal encoders. The Local Illumination-aware Fusion (LIF) module was designed to dynamically set different weights for the regions with different illumination conditions. Extensive experimental results demonstrated that M2D-LIF achieved superior performance. We believe that our framework holds potential for expansion to the other multi-modal tasks.

### v2

@@ 4.1 Datasets and Evaluation Metrics
1) DroneVehicle: This dataset [27] includes images captured by drones in urban areas under different lighting conditions. It has annotations with oriented bounding boxes for five categories: ‘car‘, ‘truck’, ‘bus’, ‘van’, and ‘freight car’. There are 28,439 pairs of RGB and IR images, with 17,990 for training, 1,469 for validation, and 8,980 for testing.
@@
Table 2: Ablation on our M2D-LIF framework with performance measured by mAP (%). The best results are highlighted in bold. Note: Params/FLOPs tested on FLIR during reference.
M2D LIF FLIR DroneVehicle LLVIP Params FLOPs
43.5 66.9 68.2 36.47M 116.7G
45.1 69.2 70.7
45.0 68.5 69.4
46.1 70.6 70.8 36.53M 123.2G
Table 3: Ablation on different ways of the weight map generation in the LIF module with performance measured by mAP (%). The best results are highlighted in bold. Sup. represents supervision.
Fusion weight FLIR DroneVehicle LLVIP
L-channel 44.2 67.4 70.2
w/o L-channel Sup. 45.0 69.8 70.3
w/ L-channel Sup. 46.1 70.6 70.8
4) Mean Average Precision (mAP): mAP is a standard metric for object detection that evaluates performance based on classification accuracy and Intersection over Union (IoU). Specifically, mAP50 denotes the mAP across all classes at a fixed IoU threshold of 0.50. The mAP is the average of mAP values calculated over a range of IoU thresholds from 0.50 to 0.95 with a step of 0.05.
@@ 4.2 Implementation Details
All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs. We employ CSPDarknet53 as the backbone network and modified it to process dual-modal inputs. The network is optimized using SGD with a momentum of 0.9370.9370.937 and a weight decay of $5\times 10^{-4}$. We initialize the learning rate at $1\times 10^{-2}$ and gradually decrease it to $1\times 10^{-4}$ throughout the training process. In the testing phase, we set the confidence threshold to 0.250.250.25 to filter detection results. If there is no special state, the ablation study results on the DroneVehicle dataset are conducted on the validation set.
4.3 Ablation Study
1) Study on Each Component: To evaluate the effectiveness of the M2D-LIF framework, we conducted ablation studies on the M2D and LIF modules. As shown in Table 2, we incrementally applied these components to the baseline model to evaluate their individual contributions. Without either of these two modules, the model achieves mAP scores of 44.4%, 66.9%, and 68.2% on the FLIR, DroneVehicle, and LLVIP datasets, respectively. The integration of the M2D training improves model performance by 1.6% in FLIR, 2.3% in DroneVehicle, and 2.5% in LLVIP and LIF fusion module improves by 1.5%, 1.6%, and 1.2%, respectively. The best results are achieved when combining both the M2D and LIF modules, yielding significant mAP increases of 2.6%, 3.7%, and 2.6% over the baseline while requiring only a marginal 0.1% parameter increase and 5.6% additional computational overhead. These improvements demonstrate the significant impact of our complete M2D-LIF framework.
@@ 4.3 Ablation Study
Figure 5: Visualization of the Fusion Degradation phenomenon across the DroneVehicle (first and second rows) and FLIR (third and fourth rows) datasets. Columns from left to right represent infrared modality, RGB modality, naive add method, weighted-based method, attention-based method, our M2D-LIF, and the ground truth. The Green boxes indicate correctly detected objects, while the red dashed boxes represent the Fusion Degradation phenomenon, where objects are detected by mono-modal models but missed by the multi-modal object detection model. The Orange boxes emphasize how our proposed method effectively resolves this issue. Zoom in for details.
Figure 6: Visualization of the Weight Map $W_{V}$ in Equation 8. From left to right are the infrared image, the RGB image, the weight map, and the grayscale color corresponding to the value.
2) Study on the LIF: To demonstrate that our proposed LIF module surpasses simply sensing illumination, we conducted a comparative experiment against the method directly using the normalized L-Channel from LAB color space as fusion weight. As shown in Table 3, our LIF module outperforms this approach by 1.9%, 3.2%, and 0.6% on the FLIR, DroneVehicle, and LLVIP datasets, respectively. Simply using the normalized L-Channel as the fusion weight cannot enable dynamic learning of the fusion weight, which is not conducive to the complementary fusion of multiple modalities. In the FLIR dataset, it even underperforms compared to the baseline method. These results further confirm that our LIF module can not only utilize the illumination context but also effectively leverage the complementary relationship between different modalities.
3) Study on the Hyper-parameter $\beta$: To investigate the impact of the $\beta$ value in the Equation 8 on M2D-LIF framework performance, we conducted ablation studies across a pre-defined set of $\beta$ values. As shown in Figure 4, The bar chart represents the AP of different classes, and the line chart represents the mAP. When beta takes the value of 0.4, our method achieves the best result. If the value of beta exceeds 0.8, it may lead to the modality imbalance, thus causing a serious decline in performance. More ablation studies are detailed in the supplementary material.
@@ 4.4 Visualization
Table 4: Camparison of the performance measured by mAP50, mAP on the DroneVehicle dataset. The best results are highlighted in red and the second-place results are highlighted in blue. Noted that it uses the ‘OBB’ detectors.
@@ 4.5 Comparison With State-of-the-Art Methods
1) On DroneVehicle: We evaluated the detection performance and parameter count of our method against four mono-modal oriented bounding box (OBB) detection methods, as well as eight SOTA multi-modal OBB detection methods on the DroneVehicle test and validation set. Table 4 presents the comparative results. Notably, our M2D-LIF framework achieves the highest mAP50 and mAP of 81.4% and 68.1% on the test set. Additionally, it recorded the highest mAP50 of 85.2% on the validation set. Furthermore, our method attained the highest mAP50 across all five categories on the validation set and three first-place and two second-place rankings on the test set. Our method has not only achieved state-of-the-art (SOTA) performance, but also the parameter count of our M2D-LIF framework is optimal with 37.1M. It has only increased by 0.1M compared to the baseline model in the test phase, which demonstrates its superiority over other multi-modal methods.
2) On FLIR and LLVIP: Furthermore, we conducted comprehensive comparisons of our method against state-of-the-art object detection methods on the FLIR and LLVIP datasets, involving four mono-modal and seven multi-modal methods, as presented in Table 5. Our M2D-LIF framework achieves 46.1% and 70.8% of mAP on the FLIR and LLVIP datasets, respectively, while only employing 36.5M parameters, the second lowest among multi-modal methods. The above results underscore the effectiveness of our M2D-LIF framework in handling diverse datasets. The superior performance on all three datasets highlights the adaptability of the proposed framework. Our method not only competes with but often surpasses existing state-of-the-art multi-modal methods while achieving higher computational efficiency with a lower parameters.
@@ 5 Conclusions
In this paper, we first observed the fusion degradation phenomenon. Through linear probing evaluation, we further identified mono-modality insufficient learning as its underlying cause of this phenomenon. To address this issue, we rethought the multi-modal object detection task from the mono-modality learning perspective and constructed an end-to-end M2D-LIF framework. The Mono-Modality Distillation (M2D) method was designed to enhance the feature extraction capability of the multi-modal encoders. The Local Illumination-aware Fusion (LIF) module was designed to dynamically set different weights for the regions with different illumination conditions. Extensive experimental results demonstrated that M2D-LIF achieved superior performance. We believe that our framework holds potential for expansion to other multi-modal tasks.
Acknowledgement
This work was supported by the Fundamental Research Funds for the Central Universities.